Insulator capacitance of MOSFET 10 micro F . width of depletion layer 0.2 cm and permitivity is 33 . calculate total capacitance


1
Expert's answer
2021-06-15T10:57:01-0400

"C_{dep}=\frac{\epsilon}{w_{dep}}\newline

\epsilon=33 \newline

w_{dep}=0.2cm=0.002m \newline

C_{in}=10\mu F \newline

C_{dep}=\frac{33} {0.002}=16500F=0.0165\mu F \newline"

Total capacitance is

"(C)=\frac{C_{dep} \times C_{in}}{C_{dep} + C_{in}} \newline

C=\frac {0.0165 \times 10}{0.0165+10} \newline

C=\frac {0.165}{10.0165} \newline

C= 0.0165\mu F=16.5nF"